You borrowed $28,000 at 11.3% interest rate to purchase a new truck.

Payments of $1,500 are made every 3 months.

What is the final payment?
Answer #1

Using financial calculator

Step 1: Number of payments required

I/Y=11.3%/4

PMT=1500

PV=-28000

FV=0

CPT N=26.89913254

So, 27 payments will be required where 26 payments will be 1500 each and last or final payment will be smaller

Step 2: Amount remaining after 26 payments

N=26

I/Y=11.3%/4

PMT=1500

PV=-28000

CPT FV=1313.48

So, final payment=1313.48*(1+11.3%/4)=1350.58581